What is the interest on Principal of $2,400, borrowed at 13.75% rate, for 5 months?
Answer by stanbon(75887)

What is the interest on Principal of $2,400, borrowed at 13.75% rate, for 5 months?
---------
I = P*r*t
-----
I = 2400*0.1375*(5/12) = $137.50
------------
Cheers,
Stan H.
----------